A car accelerates from rest at a constant rate for the first 10s and covers a distance of x. It covers a distance of y in the next 10s with the same acceleration. Which of the following is true
a)x=3y
b)y=3x
c)x=y
d)y=2x
1
Expert's answer
2018-08-17T10:10:08-0400
Let S_1 is the distance that car covers for the first 10 s
S_1=(at_1^2)/2=(a×〖10〗^2)/2=50a=x
Let S_2 is the distance for the first 20 s
S_2=(at_2^2)/2=(a×〖20〗^2)/2=200a
Thus the distance for second 10 s
S_2-S_1=200a-50a=150a=y
Finally we get
y=3x
Answer: b) y=3x